OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)

OpenStudy (jackellyn):

if we put it into y=mx+b form we can find the y intercept 2x-y=8 -y=8-2x y=-8+2x y=2x-8 so the y intercept is -8 to find the x intercept, let y=0 0=2x-8 8=2x x=4 **the opposite could've been done to find the y intercept, let x=0 y=2(0)-8 y=-8

OpenStudy (anonymous):

Thanks Jack!!